OpenStudy (anonymous):

Simplify i41 i –1 –i 1

OpenStudy (cwrw238):

i^40 = ( i^4) ^ 10 = 1 so i^41 = 1 * i = i

OpenStudy (cwrw238):

note i^4 = 1

OpenStudy (cwrw238):

the answer is i

OpenStudy (cwrw238):

remember that i^2 = -1 , i^4 = 1, i^3 = -1 * i = -1